ABC Ltd, a high-technology company, issues a $33 million IPO with an offer price of $4 per share, underwritten at $3.76 per share. The company's legal fees, ASIC registration fees, and other administrative costs are $430,000. The company's share price increases by $0.6 on the first day. What is the company's total cost of issuing the securities (in millions of dollars to three decimal places; don’t use $ sign eg $4.5766 million is 4.577)? (Remember to round the number of shares issued to a whole number)
Answer #1

Answer:

Number of share outstanding = $33,000,000 / $4.00 = 8,250,000

(i) Total underwriting cost = ($4.00 - $3.76) × 8,250,000 = $1,980,000

(ii) Legal/ASIC registration & other administrative costs = $430,000

(iii) Loss due to under-pricing on day 1 = $0.60 × 8,250,000 = $4,950,000

Total cost of issuing securities = (i) + (ii) + (iii) = $1,980,000 + $430,000 - $4,950,000

                                                                         = $7,360,000

Cost of issuing securities per share = $7,360,000 / 8,250,000 = 0.892